# Foundational GPT Model for MEG

### v1

@@ 3.1 Multi-channel Wavenet
The full Wavenet architecture can either be interpreted as forecasting with extra conditioning or as a generative encoder augmented with past brain activity. In addition, the probabilistic formulation allows converting the model into a decoder using Bayes’ rule, enabling both forecasting and decoding within the same framework:
$p(Y|X)$ $=\frac{p(X=x|Y)p(Y)}{p(X=x)}$ (13)
where $X$ is the random variable representing the data, $Y$ is the random variable representing the task labels, and $x$ is a particular sample of $X$. $p(Y)$ is the task label prior distribution which in the 118-image dataset is uniform. $p(X=x|Y)$ is the likelihood of the data given the label which we get from the above formulation of Wavenet. The only tricky part is $p(X=x)$ as this requires marginalisation over $Y$. In the case of the 118-image dataset this means that we have to run the trained model with all of the possible task labels to obtain $p(X=x)$:
$p(X=x)=\sum_{i=1}^{N}{p\left(X=x|Y=i\right)p\left(Y=i\right)}$ (14)
Thus, in a single self-supervised deep learning model we have flexibly encapsulated forecasting, encoding, and decoding, all three of the main modelling methods of M/EEG data. This unification of modelling approaches was inspired by a GitHub repository applying similar ideas to images [^3] https://github.com/cheind/autoregressive. The inverted decoder formulation also allows for iterative estimation of $p(Y|X)$ at each timestep. The author of the GitHub repository has applied this method to estimating the probability of image labels (digits 0 to 9) from pixel images, as more and more of the image was fed into the model.
3.2 Multi-channel GPT2
We set out to design a Transformer model suited for M/EEG data, while keeping the key elements that made it successful in language modelling. Specifically, we use GPT-2, a popular autoregressive Transformer variant. When adapting GPT-2 to continuous multivariate time series, the main challenges are at the input and output layers interfacing the model with the data. A particularly detailed visual description of GPT2 is given in Alammar, (2019).
Altogether a GPT2 layer is the combination of the self-attention and feedforward layers:
$\mathbf{H}^{(0)}$ $=\mathbf{X}\mathbf{W}_{e}+\mathbf{W}_{p}$ (15)
$\mathbf{Z}^{(l)}$ $=\mathrm{LN}(\mathbf{H}^{(l)}+\mathrm{MultiHeadAttention}(\mathbf{H}^{(l)}))$ (16)
$\mathbf{H}^{(l+1)}$ $=\mathrm{Dropout}(\mathrm{LN}(\mathbf{Z}^{(l)}+\mathrm{FFN}(\mathbf{Z}^{(l)})))$ (17)
$\hat{\mathbf{Y}}$ $=\mathrm{softmax}(\mathbf{H}^{(L)}\mathbf{W}^{T}_{e})$ (18)
where $\mathbf{W}_{e}\in\mathbb{R}^{Q\times E}$ embeds the discrete tokens $\mathbf{X}\in\mathbb{R}^{T\times Q}$ into $E$ dimensions. $\mathrm{LN}$ is Layer Normalisation, a regularisation technique which normalises all activations within a layer to zero mean and unit variance. $\mathbf{W}_{p}\in\mathbb{R}^{T\times E}$ contains positional encodings, providing the model with sequential order information. This is needed as GPT2 lacks recurrent or convolutional elements. Each vector in $\mathbf{W}_{p}$ indexed by $t\in(1,\dots,T)$ contains a distinct $E$-dimensional representation of position $t$. The output $\mathbf{H}^{(L)}$ is projected back to the vocabulary via the transpose embedding matrix (weight tying). Alternatively, a separate output projection can be learned. The softmax output gives a token probability distribution.
GPT-2 is trained via supervised learning to predict the next token given previous context, minimising cross-entropy loss between model outputs $\hat{\mathbf{Y}}$ and ground truth targets $\mathbf{Y}$. To enable autoregressive training, $\mathbf{Y}$ is set to $\mathbf{X}$ shifted one timestep ahead. Crucially, to prevent information leakage from future timesteps ${t+1,\dots,T}$, causal masking is applied in each self-attention layer, setting outputs that would reveal future information at position $t$ to zero.
To apply GPT2 to our continuous multichannel time series data, we take a similar approach as with Wavenet by tokenising each channel independently using the same method as before. This serves as our equivalent of the discrete set of tokens in language modelling. The same GPT2 model is applied to each channel in parallel by setting the channel dimension as the batch dimension. We call this ChannelGPT2.
The input to the model includes the position embeding as well as subject and task-stimulus embeddings. We also add a label/embedding telling GPT2 which channel the current time series is coming from:
where $+$ denotes element-wise addition, $\mathbf{X}\in\mathbb{R}^{C\times T\times Q}$ is the tokenised input, $\mathbf{W}_{c}\in\mathbb{R}^{C\times T\times E}$ are the learned channel embeddings of size $E$, which are distinct for each channel $c\in{1,\dots,C}$ but constant across time $t$. $\mathbf{Y}$ and $\mathbf{O}$ are the task and subject index matrices, mapped to their respective embeddings. As with the positional encoding $\mathbf{W}_{p}$, we simply add all embeddings (task, subject, channel) into a single representation. Note that instead of having channel-specific embeddings of the tokenised input $\mathbf{X}$ we learn the same mapping $\mathbf{W}_{e}\in\mathbb{R}^{Q\times E}$ across channels. Channel information is provided to the model through the channel embeddings.
3.3 Model interpretation
To evaluate whether Wavenet and GPT2 models accurately capture brain dynamics beyond just predictive performance, we develop several analysis techniques to interrogate what these models learn.
Data generation
Different models have distinct generation procedures. Linear AR models take Gaussian noise as input and generate one timestep at a time. Gaussian noise is added to the output, which is appended to the input sequence. This recursive process is described by:
$\mathbf{x}_{t}=\boldsymbol{\epsilon}_{t}+f({\mathbf{X}_{t-K:t-1}})$ (20)
This intuitively treats the model $f$ as a black-box infinite impulse response (IIR) filter, where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,1)$, and $K$ is the receptive field size.
For tokenised models (Wavenet and GPT2), we generate data by sampling from the predicted probability distribution and recursively feeding the sample back as input. We use top-p sampling, which samples from the ordered outputs whose cumulative probability mass exceeds $p\%$ (Holtzman et al.,, 2020). top-p sampling often works better in practice than using the full distribution by avoiding generation of low-probability tokens, and thus reducing noise.
Comparison metrics
We compare generated timeseries to real data using power spectral density (PSD), and Hidden Markov Model (HMM) statistics. We fit a separate 12-state time-domain embedding HMM (TDE-HMM) to each multivariate generated timeseries (Vidaurre et al., 2018b, ). We used the osl-dynamics package (Gohil et al.,, 2023), and set the number of embeddings to 15, the PCA projection dimensionality of the channels to 80 and the sequence length to 2000. We trained the HMMs for 20 epochs with an initial learning rate of 0.02. Once fit we compare the state timecourses of generated timeseries and real data. We also compute four summary statistics (fractional occupancy, mean lifetime, mean interval, switching rate), and compare the distributions of these statistics across the 12 states. Finally we also compare the power spectra of the timeseries corresponding to each state.
Evoked activity
For task-conditioned models, we assess reconstruction of task-dependent dynamics by feeding in task labels during generation and examining evoked responses. We compare visually average evoked responses from generated and real data, and also quantify the correlation between generated and real evoked responses. Furthermore, we epoch the state timecourses of the HMMs fit to generated and real data and compare evoked state timecourses. The variability in the evoked state timecourses is also compared.
Classification
To evaluate how well models capture task activity, we apply standard decoding models (e.g., linear classification) to generated trials and compare performance to real data. We also evaluate the generalisability of decoders trained on generated data to real data. Strong similarity in these metrics would indicate accurate modelling of task responses. For decoding both generated and real data we use a simple four-layer linear neural network introduced by (Csaky et al., 2023a, ). The first layer of the model performs a dimensionality reduction across channels.
Ablation
By removing certain model components and evaluating performance, ablation studies assess the contribution of different architectural factors. We perform ablations on conditioning and channel embeddings.
3.4 Model and training details
As our dataset of choice, we used the continuous 15-subject, 118-image data from Cichy et al., (2016). Here each subject views 118 distinct images, with 30 trials/image. For each subject, the data was bandpass filtered between 1 and 50 Hz, and a notch filter was applied to remove line noise. Subsequently, independent component analysis (ICA) artifact rejection was performed with a dimensionality of 64. Components were visually inspected for each subject, and those that exhibited clear artefactual features (e.g. eye or cardiac signals) were removed. The data was then downsampled to 100 Hz. The continuous data was split into non-overlapping validation, test, and training sets. The validation and test sets included 4 trials of each of the 118 conditions, while the training set contained the remaining 22 trials. This non-overlapping uniform splitting of the continuous data was possible due to the experimental setup during data recording.
For each model the data from each channel was tokenised independently to 256 bins using a quantisation via the mu-law algorithm discussed in Section 3.1. To achieve uniform quantisation, we first standardised each continuous-data channel, clipped values higher than 4 or lower than -4, applied per-channel maximum absolute scaling to map the data to the range (-1, 1), and finally applied the mu-law transform and 8-bit quantisation.
Our aim was to evaluate several models and methods on this dataset. Due to computational constraints and limited iteration speed over experiments and methods, all experiments were performed on a single representative subject, except in Section 2.5 where we explore our GPT2 model on all 15 subjects, and the classification and transfer learning analyses in Section 2.6 and Section 2.7.
We trained univariate AR(255) models on each channel. The AR order was chosen to match the receptive field of the other models. We implemented and trained the AR model using a single linear convolutional neural network layer. Note that we did also assess multivariate AR models (results not shown), but this did not improve performance compared to the univariate AR. We trained WavenetFullChannel with a matched receptive field of 255, two stacks of dilation blocks (7 layers per block, doubling dilation factors), 256 hidden channels, 1024 skip channels, no dropout, and a 20-dimensional task embedding. WavenetFullChannelMix had the same architecture but 128 hidden channels and 512 skip channels. We used early stopping on the validation set. This means that we ran training until overfitting was observed, and then analysed the model version with the lowest validation loss. All our analyses were performed on the distinct test set.
Our Channel-independent GPT2 (ChannelGPT2) had a variable receptive field between 128 and 256. This means that during training the model encountered examples that had a sequence length between 128 and 256, rather than all examples having the same length. GPT2 is normally trained to output all timesteps in a sequence of length $T$, given previous timesteps. However, this means that for the second timestep, the receptive field is only 1. Ideally, we wanted to match the training setup of our Wavenet models, where the receptive field is always 256. However, this would significantly slow down training as the whole forward and backward pass must be recomputed at each timestep. We opted for a trade-off, where we set the minimum receptive field to 128, ensuring efficient training and that the model is not trained to predict shorter sequence lengths.
The embedding size of all inputs (token vocabulary, position, task, channel) was set to 96, and we used 12 GPT2 layers, with 12 attention heads. We used Huggingface’s implementation [^4] https://github.com/huggingface, so the rest of the parameters were the same as in their configuration. Dropout was set to 0 and we used early stopping on the validation set.
On average the mu-law quantisation achieved low reconstruction error. We tested the reconstructed data by performing evoked analysis, and classification of the task responses, and achieved comparable performance to the raw data (results not shown). Thus, both types of tokenisation add negligible quality loss to the data.
For group trainings we used the same hyperparameters as for the single-subject trainings, except for increasing the embedding size to 240. ChanelGPT2-group proved difficult to overfit, meaning that using more data acted as a regulariser, and we stopped training when validation losses did not improve for 5 consecutive epochs.
4 Discussion
We have presented our initial efforts at developing a general forecasting model for M/EEG data. After carefully evaluating the trade-offs between various modelling approaches, we settled on two main architectures: one based on Wavenet (van den Oord et al.,, 2016), and one based on GPT-2 (Radford et al.,, 2019). These models have proven successful in the audio and natural language domains, which share similarities with the time series nature of brain signals. We systematically compared different variants of our proposed models on a visual-task MEG dataset.
We found that the forecasting performance was comparable between Wavenet and AR models according to next-timestep prediction metrics. This suggests such metrics may be limited in their ability to effectively evaluate model dynamics beyond one-step prediction. Perhaps looking at these metrics when recursively generating multiple timesteps in the future might be more informative. Generated data analysis provided more discerning model comparisons. While the channel-independent AR and Wavenet models accurately reproduced the overall power spectral density, only the Transformer-based models captured more abstract multivariate statistics like HMM state dynamics.
Critically, the ChannelGPT2 model-generated data closely matched real MEG recordings across both temporal and spectral domains. Analysis of the discovered latent brain states showed ChannelGPT2 reproduced variable oscillatory states similar to those inferred from human recordings (Vidaurre et al., 2018a, ). Each state captured distinct spectral content, while the linear and Wavenet-based models failed to achieve this degree of heterogeneity in their dynamics. It is possible that this does not indicate a failing of the Wavenet architecture, but rather that different conditioning methods may be needed. One such approach that we have not tested is using the same type of channel embeddings as for ChannelGPT2.
Multiple analyses consistently demonstrated ChannelGPT2’s strengths in realistic conditional timeseries generation. ChannelGPT2-generated evoked responses had high correlation to real MEG data. However, modelling single-trial variability and between-subject differences remain challenging areas needing further work. Scaling to multiple subjects showed promise. The model was able to adapt its generated data based on the input subject label and generate task trials with variability more similar to real recordings than a single-subject model.
Ablation studies quantified the importance of channel embeddings and task conditioning for accurate MEG modelling. Removing channel embeddings resulted in near identical generation across sensors, failing to capture spatial heterogeneity. Analysis of ChannelGPT2’s channel embeddings revealed spatial relationships between sensors were learned, with proximal channels having more similar embeddings. With incorrect or with no task labels, ChannelGPT2 produced noisy evoked responses, indicating the model leverages both timing and label semantics. Furthermore, the model trained on 0.5s trials only, was able to produce reasonable responses to longer or shorter trials, showcasing generalisation. These results demonstrates the value of multi-faceted conditioning for realistic brain data modelling.
A key investigation involved analysing the classification of generated data according to the condition labels. The trials generated by the group-level model were classified with much higher accuracy (closer to real data) than those of the single-subject model. We also demonstrated that generated data can improve decoding of real trials via transfer learning (Torrey and Shavlik,, 2010), with benefits scaling with generated data quantity. The classifiability of generated trials and transfer learning results highlight the utility of forecasting models like ChannelGPT2 for decoding real MEG data. Further analysis could involve permutation feature importance (Altmann et al.,, 2010) of the decoding model trained on generated data to gain insights into learned representations.
4.1 Comparison with previous works
The Deep Recurrent Encoder (DRE) proposed by Chehab et al., (2022) is a highly relevant architecture to our approaches, as it demonstrates the advantages of modelling spatiotemporal dynamics for encoding neural data. DRE aims to predict MEG brain responses to visual word stimuli. While motivated as an encoding model, DRE can also be viewed through the lens of forecasting, with the addition of auxiliary task features. Forecasting holds inherent advantages over pure encoding, as it enables reconstructing real data and modelling complex spatiotemporal relationships, beyond just learning abstract representations.
Banville et al., (2021) investigate three SSL tasks for learning from unlabelled EEG recordings. Each task is trained via a contrastive loss function, where the model learns to pull positive pair examples closer in a representation space while pushing negative pairs apart. They demonstrate that the representations learned via SSL on unlabelled EEG data transfer well to supervised downstream tasks, consistently improving over limited label training and matching full supervision performance.
Building on this Kostas et al., (2021) propose combining self-supervised contrastive learning with Transformer networks to enable pre-training on large amounts of unlabelled EEG data. Their approach, BErt-inspired Neural Data Representations (BENDR), uses a Transformer encoder architecture applied to learned representations of raw EEG segments. First, a temporal convolutional network extracts initial representations of the EEG time series, referred to as BENDR features. Next, a Transformer encoder module takes the BENDR features as input. Contiguous segments of the BENDR representations are randomly masked, and the model is trained via a contrastive loss to predict the original features. Fine-tuning the pretrained model significantly improves performance on supervised EEG analysis tasks compared to training just on the downstream datasets.
In a similar vein, Wang et al., (2023) propose BrainBERT, a Transformer model tailored for self-supervised pretraining on unlabelled intracranial field potential recordings. BrainBERT is trained to reconstruct randomly masked patches of time-frequency spectrograms computed from the raw voltage measurements. Compared to BENDR, BrainBERT works directly on spectrogram representations rather than learned features. Similar to BENDR, the pretrained BrainBERT model serves as a general purpose feature extractor. By training linear classifiers on top of BrainBERT embeddings, Wang et al., (2023) achieve large performance gains on neural decoding tasks compared to raw or hand-engineered input features. Critically, BrainBERT generalises very well to new subjects and electrode locations.
In the image domain, tokenisation is often abandoned, and a linear projection directly maps image patches to continuous vector representations (Dosovitskiy et al.,, 2020). Similarly, Nie et al., (2022) have designed a channel-independent Transformer architecture applied to overlapping patches of continuous time series for forecasting. While this facilitates the input, without tokens categorical outputs cannot be generated. We posit that maintaining operations over tokens and categorical outputs are desirable GPT2 features for M/EEG data. This is because we would like to output probability distributions and train using the cross-entropy loss. Previous applications of Transformer models to M/EEG data do not enforce a discrete tokenisation of the input, nor are they capable to generate new data auto-regressively. The latter allows for better interpretability of learned dynamics and artificial data simulation.
The tokenisation can happen either before or after mixing information across channels. The latter matches GPT2’s original design. One example of this is vector quantisation, which is used to tokenise multiple channels in Jukebox, a successful autoregressive Transformer model used on audio data (Dhariwal et al.,, 2020). Before training the Transformer, a hierarchical VQ-VAE (vector quantized variational autoencoder (Van Den Oord et al.,, 2017)) learns discrete codes (tokens) from raw audio. Once trained, VQ-VAE can map a continuous time series to a discrete token sequence $\mathbf{z}$. In the second step of Jukebox, the VQ-VAE is kept fixed, and the discrete tokens are used to learn an autoregressive Transformer.
Importantly, VQ-VAE is applied to single-channel audio to compress the temporal dimension into discrete codes. As our aim is to mix information across the channels we would primarily want to apply vector quantisation to the channel dimension, to have a discrete token at each timestep, or perhaps across a few timesteps. While an adaptation of this could work on MEG data, we experimented with a simpler non-deep learning method. We tried applying vector tokenisation on small groups of channels using the Residual Quantiser algorithm (Babenko and Lempitsky,, 2014). Unfortunately, this approach to include information from all channels in the input sequence resulted in worse generation capabilities. See Supplementary Section A.1 for details on this method, and Supplementary Section A.3 and Section A.4 for results.
4.2 Limitations and future work
The full potential of self-supervised learning is only realised with large-scale data. This remains challenging for brain imaging compared to vision and language. Lowering barriers to data access and promoting data sharing is critical to realise the promise of foundation models in neuroimaging (Poldrack and Gorgolewski,, 2014).
A core limitation of the channel-independent GPT2 model is no direct leveraging of cross-channel information for each sensor prediction. Our approach leveraging vector quantisation performed worse. We think that maintaining the innate inductive biases of Transformers, which emphasise 1D sequence modelling on embeddings of discrete tokens, is paramount. Different architectures or more data may enable proper utilisation of cross-channel dependencies. We tried various other approaches to mixing channel information beyond those reported, without success. For the Wavenet model, we incorporated all channels in the input by concatenating embeddings, and for the GPT2 models, we tried mixing channels with convolutions. We tried concatenating the output of each channel and then predicting from this shared output using a different projection for each channel. We also attempted to increase receptive field, dropout, and model size. One limitation in our approaches is the use of a next-timestep prediction loss. Future work should continue exploring architectures and different self-supervised or multi-timestep losses to leverage cross-channel information and improve modelling capabilities.
Some of our findings substantiated that predicting the next timestep may not serve as a robust measure of model performance. Future research should contemplate adopting multi-timestep or contrastive loss frameworks. A plausible strategy could involve deploying the VQ-VAE model across both channel and temporal dimensions, aiming to distill a more coarse sequence of discrete tokens. Nevertheless, any quantisation-centric approach must carefully consider reconstruction error. We posit that a significant portion of the signal dynamics should be entrusted to the Transformer, given its adeptness in capturing complex dynamics.
A constraint in our modelling approach is its reliance on categorical task stimuli labels. Such an approach, while effective in our context, does not readily lend itself to scalability across diverse tasks and datasets. However, it is conceivable to construct robust representations tailored for various stimulus modalities—ranging from images to audio. These representations can then serve as conditioning embeddings. As shown by Défossez et al., (2022), tools such as wav2vec (Baevski et al.,, 2020) can be leveraged to derive informative representations of auditory stimuli.
Transfer learning also requires more thorough evaluation across diverse decoding tasks. It would be important to also investigate other more direct finetuning or transfer learning approaches of the forecasting model akin to vision or language domains. These could involve additional output layers and losses for finetuning on downstream tasks.
In conclusion, this work demonstrates that deep forecasting models can accurately reproduce complex neural dynamics of both ongoing and task-related activity and provides an extensive analysis methodology. Future work should explore more flexible conditioning, study different self-supervised and transfer learning frameworks, and critically, apply similar analyses when scaling up across diverse, large electrophysiology datasets. This has the potential to enable powerful transfer learning and advance foundational brain modelling and decoding.
Acknowledgments
Appendix A Supplementary Material
A.1 FlatGPT2
FlatGPT2 is motivated by the observation that language models include extra information such as context within the sequence, instead of the feature space. We provide the detailed procedure below.
Directly vector quantising 300 channels to any vocabulary size would result in poor reconstruction. In FlatGPT2 we perform the tokenisation on small groups of channels instead. First, we compute the covariance over channels in the training data. Then, we apply K-means clustering (Hartigan and Wong,, 1979) on the covariance matrix to group channels into buckets. This ensures that each bucket contains channels with high covariance. This is important because tokenising a feature space (group of channels) with high covariance can be done with fewer tokens while maintaining low reconstruction error. We set the number of clusters ($B=30$) based on manual tuning on the training data. Each cluster/bucket can contain a variable number of channels, usually between 5 and 20.
After assigning channels to buckets we apply the Residual Quantiser algorithm (Babenko and Lempitsky,, 2014) from the faiss library [^5] https://github.com/facebookresearch/faiss/wiki/Additive-quantizers to each bucket $b$ separately. This is a powerful additive quantiser (Liu et al.,, 2015) that achieves good reconstruction error with a relatively small vocabulary size $V$. Note that the total number of tokens, i.e. the vocabulary size will be $BV$, since we have $B$ quantisers. Once fit to the training data the quantiser is fixed and can be applied to new data.
Mathematically, the covariance is obtained by:
$\forall i,j\in{1,\dots,C}\quad\mathbf{C}_{ij}=\frac{1}{T}\sum_{t=1}^{T}(x_{t,i}-\mu_{i})(x_{t,j}-\mu_{j})$ (21)
Where $x_{t,i}$ is the $i^{th}$ channel at timestep $t$, $\mu_{i}$ is the mean of channel $i$ over all timesteps, and $C$ is the total number of channels. $\mathbf{C}$ is a symmetric matrix, and thus the feature and variable dimensions of K-means are the same. K-means computes buckets $\mathcal{C}_{1},\dots,\mathcal{C}_{B}$ which partition channels $C$ into distinct sets with high within-bucket covariance.
The residual quantiser $Q_{b}$ learns a codebook $\mathbf{C}_{b}\in\mathbb{R}^{V\times|\mathcal{C}_{b}|}$ for each bucket $\mathcal{C}_{b}$:
$\forall t\in{1,\dots,T}\quad z_{t,b}=Q_{b}(\mathbf{x}_{t,b};\mathbf{C}_{b})$ (22)
Where $z_{t,b}$ is the quantised representation (token/code) at timestep $t$ of the channels $\mathbf{x}_{t,b}\in\mathbb{R}^{|\mathcal{C}_{b}|}$ in $\mathcal{C}_{b}$. The encoding in the quantiser is sequential, thus at stage $m$ of the encoding of $\mathbf{x}_{t,b}$, the quantiser picks the entry $i_{m}$ that best reconstructs the residual of $\mathbf{x}_{t,b}$ w.r.t. the previous encoding steps:
[…]
where $\mathbf{T}_{m}$ is a table of size $K_{m}$ containing $|\mathcal{C}_{b}|$ dimensional vectors. For notational simplicity we omit the index $b$ from $i_{m}$ and $\mathbf{T}_{m}$ in the above. The quantisation provides a vector $[i_{1},...,i_{M}]$, where each element $i_{m}$ comes from a set of size $\lceil\log_{2}(K_{m})\rceil$ bits. This bit vector representation can be easily transformed to token indices ranging from 1 to $V=\sum_{m=1}^{M}\lceil\log_{2}(K_{m})\rceil$. Note that this table description of the discrete code is just a different representation of the overall codebook $\mathbf{C}_{b}$. A code $[i_{1},...,i_{M}]$ can be reconstructed to obtain $\hat{\mathbf{x}}_{t,b}$ by retrieving the corresponding vectors from $(\mathbf{T}_{1},\dots,\mathbf{T}_{m})$ and adding them up. Reconstruction error is computed by comparing $\hat{\mathbf{x}}_{t}$ and $\mathbf{x}_{t}$.
By using 30 buckets we obtain 30 tokens per timestep, which is already a 10-fold reduction of the original dimension space, but we have not reached our initial goal of having 1 token per timestep. To achieve this, we flatten the feature dimension (buckets) when feeding tokens to GPT2, hence the name FlatGPT2. Our total sequence length then becomes $B\cdot T$, where $B$ is the number of buckets and $T$ is the number of timesteps. This approach is also motivated by the observation that language models include extra information such as context within the sequence, instead of the feature space. Thus, when predicting the token of bucket $b$, we treat the previous timesteps of the other buckets as contextual information.
We also add an extra separator token $z_{sep}$ between sequences of buckets corresponding to the same timestep to facilitate distinction between the bucket and time dimensions. An input sequence to FlatGPT2 consists of tokens $z_{t,b}$ following a fixed order:
$\mathbf{z}=($ $z_{sep},z_{t=1,b=1},z_{t=1,b=2},\dots,z_{t=1,b=B},$ (24)
$z_{sep},z_{t=2,b=1},z_{t=2,b=2},\dots,z_{t=2,b=B},$ (25)
$z_{sep},\dots,\dots,z_{t=T,b=B})$ (26)
For each codebook $\mathbf{C}_{b}$ a separate embedding $\mathbf{W}_{e,b}\in\mathbb{R}^{V\times E}$ is learned. As in ChannelGPT2 we add the appropriate conditioning embeddings to the input embedding with appropriate flattening across the channel/bucket dimension:
$\mathbf{H}^{(0)}$ $=\mathbf{Z}\mathbf{W}_{e}+\mathbf{W}_{p}+\mathbf{Y}\mathbf{W}_{y}+\mathbf{O}\mathbf{W}_{o}+\mathbf{W}_{c}+\mathbf{W}_{t}$ (27)
where $+$ denotes element-wise addition and $\mathbf{Z}\in\mathbb{R}^{(B+1)T\times V}$ is the one-hot version of $\mathbf{z}$. The task labels $\mathbf{Y}$ can vary across time, but are the same across the buckets of one timepoint. $\mathbf{W}_{c}$ now contains distinct embeddings of buckets $b\in(1,\dots,B)$, which are the same across timesteps. We also augment the input with $\mathbf{W}_{t}$, containing distinct embeddings for timesteps $t\in(1,\dots,T)$, which are the same across buckets. This is the timestep version of $\mathbf{W}_{c}$.
As usual, the model is trained to autoregressively predict the next token in the sequence given all previous inputs. At timestep $t$ and bucket $b$ the model has access to the tokens $\mathbf{z}_{1:t-1}$ from all buckets (and thus information from all channels), and the tokens $\mathbf{z}_{t,1:b}$, and has to predict token $z_{t,b+1}$. The buckets of the same timestep are predicted sequentially, thus, bucket ordering could influence results. We use an arbitrary bucket ordering and do not experiment with different orderings of the input sequence.
Note that at the last bucket $B$ in each timestep the prediction should be token $z_{sep}$, however, we simply discard this prediction during loss computation, as we do not require the model to predict separator tokens. The structure of the sequence already constrains the predictions such that a new timestep begins after every $B$ tokens. Conversely, when computing the prediction at input token $z_{sep}$, the target is the token with bucket $b=1$ of the next timestep. This is useful as in theory we could start the recursive generation of data with a single $z_{sep}$ token.
[…]

### v2

@@ A.1 Multi-channel Wavenet
A.2 Evaluation details
We fit 12-state time-domain embedding HMMs using osl-dynamics (Gohil et al., 2023) with 15 embeddings and PCA to 80 dimensions (sequence length 2000). Decoding uses the 4-layer linear network of Csaky et al. (2023b).
A.3 Preprocessing and data splits
We bandpass filter 1–50 Hz, apply a notch filter for line noise, perform ICA artifact rejection (64 components with manual component rejection), and downsample to 100 Hz. We split the continuous recording into non-overlapping blocks corresponding to trials, holding out 4 trials/condition for validation and 4 for test.
A.4 Tokenization quality
We verified that de-quantized signals preserve evoked responses and decoding performance compared to raw continuous data; reconstruction error is low (less than 5%) for both $\mu$-law and linear quantization.
A.5 Hyperparameters
Model Univariate Tokenised Linear
AR(255) yes no yes
WFC yes yes no
WFCM no yes no
GPT2MEG yes yes no
Table 1: Summary of evaluated models. “Univariate” treats channels as a batch dimension during training; “Linear” denotes linear dynamics in the model class.
We match receptive fields across deep models (255 samples). WaveNet variants use two dilation stacks (7 layers each) with early stopping. We set dilation and residual channels to 128, and skip channels to 512. GPT2MEG uses 12 layers and 12 attention heads with embedding size 96 for single-subject models and 240 for group models; optimization uses Adam (Kingma and Ba, 2015) with early stopping. Batch size is set to the number of channels, so 1 full example for channel-independent models. Please see Table 1 for a summary on model variants.
(a) WavenetFullChannelMix p=0.72
(b) WavenetFullChannelMix p=0.80
Figure 8: PSDs for WavenetFullChannelMix at different top-$p$ values; (a) $p=0.72$ and (b) $p=0.80$ for a model trained on a single subject. Each line represents a different MEG sensor/channel.
A.6 State-specific PSDs of HMM states
In addition to state statistics, we also computed the power spectra of each state across the timeseries. In real MEG data, different HMM states can capture oscillatory activity with specific frequencies (Vidaurre et al., 2018b). The extracted power spectra from the different inferred state time courses are shown in Figure 9. We can see that the HMM trained on the real MEG data contains many states that capture the 10 Hz peak, with fewer states having a 20 Hz peak. It is also clear that the states of the HMM fitted to the WavenetFullChannelMix generated timeseries do not contain these spectral peaks. While the AR(255) does contain states with a 10 Hz peak, the shape does not match the data well, and also states do not show the same variability as in real data.
In contrast, GPT2MEG, matches the state PSDs of the real data very well, further demonstrating the superiority of Transformer models in capturing complex neural dynamics. While WavenetFullChannel also improves substantially over the AR(255) power spectra, it falls short in capturing the 20 Hz peak and the heterogeneity between states observed in the real data and the generated data of GPT2MEG. While we have not conducted this specific analysis on multiple subjects, we see no reason why the main findings would not generalize.
(a) Data
(b) AR(255)
(c) WFCM
(d) WFC
(e) GPT2MEG
Figure 9: Power spectral density of HMM states, for an HMM inferred on (a) real MEG data from a single sample subject, and (b)-(e) data generated from the different forecasting models trained on the single sample subject. WFCM refers to WavenetFullChannelMix. Each line is the PSD of a different state. Note that states are not matched across models. Horizontal axis represents frequency in Hz.
A.7 Additional evoked-response analyses
(a) WFC mean
(b) WFCM mean
(c) GPT2MEG mean
(d) WFC var.
(e) WFCM var.
(f) GPT2MEG var.
Figure 10: (a)-(c) Correlation between the time-courses of the mean (over individual epochs) evoked responses; from the real MEG data for a single sample subject and the mean evoked responses from data generated by the different forecasting models trained on the single sample subject. (d)-(e) Correlation between the time-courses of the variance (over individual epochs) of the mean evoked responses; from the real MEG data for a single sample subject and the mean evoked responses from data generated by the different forecasting models trained on the single sample subject. For all figures, the correlation values are visualised across sensors. WFC refers to WavenetFullChannel and WFCM refers to WavenetFullChannelMix. Darker reds indicate higher correlation.
Figure 11: Evoked response state timecourses of HMMs inferred on a single sample subject; using real MEG data (top), and on generated data from each of our task-conditioned models trained on the single sample subject. Note that the HMM states are not matched between models. Image presentation starts at 0 seconds and ends at 0.5 seconds. ChannelGPT refers to GPT2MEG.
As expected, GPT2MEG generates data with evoked responses that have much higher correlation with evoked responses from real data, and slightly higher correlation in visual areas compared to other channels, matching the known topography of visual evoked responses. In other models the correlation is low, and spatially better in frontal areas, likely because the evoked responses here are noisier providing an easier fit.
Figure 10 also shows the correlation between the variance (over individual epochs) time-courses of the mean evoked response obtained from the actual data and the evoked responses obtained from data generated by each model. This captures a measure of the ability of the models to represent the trial-to-trial variability found in the real data. Again, GPT2MEG generates data that has the highest correlations with the real data, with higher values in channels in the back of the head, appropriately capturing the topography of response variability. Other models have similar spatial distribution, and notably WavenetFullChannel also produces evoked responses with variance partially matching the real data.
Finally, a different way to assess task-related activity is to examine the evoked state time-courses from the HMMs fitted on the real and model generated timeseries. Rather than looking at individual channels, this provides an overall view of which HMM state gets activated when, during individual trials. This is computed by simply epoching the state timecourse, and averaging over all trials. We plot these for the real data and each generated timeseries in Figure 11. As expected, the HMM trained on models other than GPT2MEG shows poor evoked state time-courses. GPT2MEG generated data produces states with similar evoked dynamics and variability as the real data. In the next section we show how this generalizes over multiple subjects.
A.8 Additional ablations and channel-embedding analysis
We also investigated the contributions of the channel and condition embeddings, by training two separate ablated models. As shown in Figure 12, removing the channel embeddings resulted in very similar PSD across channels in the generated data, indicating that the model relies heavily on these embeddings to adapt generation per channel. The evoked responses in Figure 13 confirm that without channel embeddings, variability between channels is reduced. Removing the condition embeddings resulted in noisier power spectra of the generated data and no 20 Hz peak.
(a) Full GPT2MEG
(b) No channel embedding
(c) No condition embedding
Figure 12: Comparison of generated power spectra with different ablations: (a) full GPT2MEG model, (b) GPT2MEG with ablated channel embeddings and (c) GPT2MEG with ablated condition embeddings. Shown for a single sample subject. Both channel and condition embeddings are critical for accurate spectral content.
(a) Channel 1 (MEG1011)
(b) Channel 2 (MEG1813)
(c) Channel 3 (MEG2332)
Figure 13: Comparison of generated evoked responses with ablated channel embeddings in the GPT2MEG model, shown across 3 representative channels (a)-(c) and for a single sample subject. Without channel embeddings the model fails to adapt evoked responses to different channels. The stimulus onset is at 0 seconds and the offset is at 0.5 seconds. ChannelGPT refers to GPT2MEG.
(a) PCA
(b) Correlation
Figure 14: Visualisation of channel embeddings. (a) 2D projection of the channel embeddings from GPT2MEG-group with PCA. Channels are coloured by their location on the scalp grouped into 5 major brain areas. (b) Plotting pairwise Euclidean distances of channels in real, physical space versus embedding space. Sensors that are near to each other in the real sensor montage tend to have more similar embeddings. Each point represents a different pair of channels. Correlation is 0.45.
(a) 1st component
(b) 2nd component
(c) 3rd component
[…]